hi,friends.. Q.if the point P(x,y) be equidistant from points A(a+b,a-b),B(a-b,a+b) then prove that bx=ay. (1)geometric (2)vectorial methods

hi,friends..
Q.if the point P(x,y) be equidistant from points A(a+b,a-b),B(a-b,a+b) then prove that bx=ay.
(1)geometric (2)vectorial methods

Grade:12th Pass

1 Answers

Chetan Mandayam Nayakar
13 years ago

(x-a-b)^2 + (y-a+b)^2 =(x-a+b)^2 +(y-a-b)^2, upon simplification bx=ay